OpenStudy (anonymous):

Find the probability that none of the three bulbs in a traffic signal must be replaced during the first 2700 hours of operation if the probability that a bulb must be replaced is a random variable b with density f(b) = 6(0.25 – (b - 2.5)2) when 2 < b < 3 and 0 otherwise, where b is time measured in multiples of 1000 hours. Find the probability that none of the three bulbs in a traffic signal must be replaced during the first 2700 hours of operation if the probability that a bulb must be replaced is a random variable b with density f(b) = 6(0.25 – (b - 2.5)2) when 2 < b < 3 and 0 otherwise, where b is time measured in multiples of 1000 hours. @Mathematics

OpenStudy (phi):

1-f(2.7)^3

OpenStudy (zarkon):

there should be an integral in your solution

OpenStudy (phi):

OK, it's a density... so integrate 0 to 2.7 f(b) to get the probability of failure for one bulb in the first 2700 hours?

OpenStudy (zarkon):

2 to 2.7

OpenStudy (anonymous):

yeah, i have an integral but is the range from 2 to 2.7. that is what i got. i don't understand how it should be from 0 to 2.7

OpenStudy (phi):

Yep, I meant 2...

OpenStudy (anonymous):

the answer i got from the integral is 183.184. is that correct?

OpenStudy (zarkon):

no...it is a density...it must be <= 1

OpenStudy (anonymous):

okay. heres what i did. I simplified the equation f(b) to be equal to -6b^2 +30b-36 and integrated that with the range from 2 to 2.7. Is that the right equation to use or am i doing it wrong

OpenStudy (zarkon):

do u-substitution

OpenStudy (zarkon):

I'll be back on later(need to go to my office)...if you haven't received a good answer I'll help you.

OpenStudy (anonymous):

okay. thank you anyway for your help so far

OpenStudy (phi):

After you integrate, you get -2b^3 + 5b^2- 36b evaluated between 2 and 2.7 this gives 0.784 so, back to the problem, prob of no failures is 1-0.784^3

OpenStudy (zarkon):

P(no failures)=P(red doesn't fail and yellow doesn't fail and green doesn't fail) if we assume independence =P(red doesn't fail)P(yellow doesn't fail)P(green doesn't fail) let \[p=\int\limits_{2.7}^{3}f(b)db\] then P(red doesn't fail)P(yellow doesn't fail)P(green doesn't fail) \[=p\times p\times p=p^3\] or let \[q=\int\limits_{2}^{2.7}f(b)db\] then the prob of not failing is 1-q then P(red doesn't fail)P(yellow doesn't fail)P(green doesn't fail) \[=(1-q)\times(1-q)\times(1-q)=(1-q)^3\] =.010077696 in either case

OpenStudy (anonymous):

what is the value for q though after integration? I keep getting big numbers despite u substitution

OpenStudy (zarkon):

.216

OpenStudy (zarkon):

is 1-q

OpenStudy (phi):

I see I am out to lunch this morning. But at least I can answer how to evaluate the integral: \[\int\limits_{2}^{2.7}-6b^2 +30b-36 db= -2b^3+15b^2-36b \mid_2 ^{2.7}=\] \[-2(2.7)^3+15(2.7)^2-36(2.7) - (-2(2)^3 +15(2)^2-36(2))\]